If (x,y) = (4.8, B) gives the coordinates of a point on the graph of
y = 5/8 x + 5 then what is the value of B?

Answer by nerdybill(7384)

If (x,y) = (4.8, B) gives the coordinates of a point on the graph of
y = 5/8 x + 5 then what is the value of B?
.
The problem gives x=4.8 -- plug this into:
y = 5/8 x + 5
y = 5/8(4.8) + 5
y = 5(0.6) + 5
y = 3 + 5
y = 8
.
Therefore,
B = 8